OpenStudy (anonymous):

The sum of two numbers is 45 and the difference is 23. What are the numbers?

OpenStudy (anonymous):

x + y = 45 x - y = 23 ----------add 2x = 68 x = 34 x + y = 45 34 + y = 45 y = 45 - 34 y = 11

Directrix (directrix):

Let x be one number The second is x - 23 x + x -23 = 45 2x - 23 = 45 Solve for x. Subtract 23 to get the other number.

OpenStudy (jahnavi):

the numbers are 34 and 11, as 34+11=45 and 34-11= 23

OpenStudy (anonymous):

omg thank you ! this actually helped!